OpenStudy (anonymous):

can someone please help me!! Explain whether the polynomials below are written in standard form. If not, use complete sentences to explain how to write the polynomial in standard form. 1.3a3b – 4ab2 + a2 2.–8x2y2 + 10xy4 – 6x2y

OpenStudy (anonymous):

well those with the sum of powers being high are written first

OpenStudy (anonymous):

so lol what does that mean? i'm so confused

OpenStudy (anonymous):

a3b what is the sum of powers here?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

no a3b = a3b1 sum of powers = 3+1 = 4

OpenStudy (anonymous):

oh my okay...so is that how to solve the problem?

OpenStudy (anonymous):

yea find it for the rest to and arrange them in decreasiong order of their 'sum of powers'

OpenStudy (anonymous):

could you walk me through the first one ? please

OpenStudy (anonymous):

3a3b – 4ab2 + a2 a3b = 3+1 = 4 ab2 = 1+2 = 3 a2 = 2 already in decreasing order 4,3,2 so this is standard form

OpenStudy (anonymous):

ohhhhhhhhhhhh okay! lol thanks

OpenStudy (anonymous):

hahaa u r welcome

OpenStudy (anonymous):

u may close the question now lol

OpenStudy (anonymous):

2. –8x2y2 + 10xy4 – 6x2y x2y2= 2+2=4 xy4=4+1=5 xy2=1+2=3 Not in decending order is this correct?

OpenStudy (anonymous):

awsumm, yepp

OpenStudy (anonymous):

alrightyy thanks again :)

OpenStudy (anonymous):

welcome again lol :D